OpenStudy (error1603):

How do you solve 8√5+3√5-√5 I guessed 11 but im not sure if they I went about it using like terms is right? for instance I added 8√5+3√5 which I got 11√5 then subtracted ↓5 from it which I got 11

OpenStudy (error1603):

yeah

OpenStudy (error1603):

so

OpenStudy (pawanyadav):

OK tell me what is 11x-x

OpenStudy (pawanyadav):

Is it 11 or 10x.

OpenStudy (error1603):

11

OpenStudy (error1603):

So 11√5 - √5=10√5?

OpenStudy (pawanyadav):

OK try this 2x - x= x(2-1)=x(1)=x Can you understand it

OpenStudy (pawanyadav):

Yes, you can't cancelled √5 from it

OpenStudy (pawanyadav):

Got it, it is 10√5

OpenStudy (error1603):

Yeah I did it this way 8 apples + 3 apples - 1 apple = 10 apples 8 √5 + 3√5 - √5 = 10 √5

OpenStudy (pawanyadav):

Good